(2) It is desired to deliver 60 mº/hr of water at 20°C through horizontal asphalted cast iron pipe. Estimate the pipe diametePLEASE USE MOODY CHART WITH PICTURES! Thanks in advance!
Answer #1

Ques) AP = hf 40 x103 998x9.81 FLV? ogo 4:0856 = fx100 V2 - 2x9:81xd Q=AV Geno) = (302) v V= ( 902129) from egocis 40856 = f*0.038 - =1780.07 d=0.11637m a: Re-20177.56 = 181972.8587 0:11637 from moody chart = 0.0012 = 0.01031 0.11637 f= 0.0386 our as